# PS-WL: A Probability-Sensitive Wear Leveling scheme for SSD array scaling

### v1

@@ 3.4 Wear Leveling Control
$\frac{|L_{o}(t)-L_{s}(t)|}{L_{o}(t)}<\theta_{exit}$ (8)
where $\theta_{exit}$ is a preset exit threshold.
@@ 4.1 Methodology
To evaluate the performance of our proposed PS-WL scheme, we conduct a comprehensive comparison against three baseline approaches: SWANS[8], Lazy-WL[9], and a custom ablation study variant, DegradedWL, with different scaling methods in various RAID configurations. We used the Alibaba Cloud Block Storage Traces[19], a write-intensive, large-scale I/O trace for our experiments. The experiment is conducted five times in order to calculate the average value of each metric and reduce the impact of random variation.
  • PS-WL: Our proposed Probability-Sensing Wear Leveling scheme, which uses a dynamic hotness vector and PID control to manage data migration based on real-time disk failure probability.
  • SWANS & Lazy-WL: Two representative state-of-the-art wear-leveling schemes used as performance benchmarks.
  • DegradedWL: To isolate and validate the effectiveness of the probability-sensing component, we designed this crucial ablation study variant. DegradedWL employs the identical data protection area and PID control logic as PS-WL but reverts to using the simpler P/E cycle count as its lifetime indicator, rather than the dynamic failure probability. This allows for a direct comparison of the intelligence provided by our probability-sensing model.
@@ 4.2 Experimental Results
We conducted tests on the SSD emulator we designed for (k,k’) scaling scenario where k is the number of original disk and k’ is the number of scaling disk. To quantify the direct benefit of our probability-sensing model, we designed a specific experiment comparing PS-WL against other WL approach. The simulation is configured to initiate the wear-leveling scheme at a critical juncture: when the highest failure probability of any single disk in the array first reaches 0.01% .
@@ 4.2.1 Average Response Time
A shorter average response time reflects a lower I/O latency and thus better system performance. As shown in Fig. 1 , PS-WL demonstrates a dramatic reduction in latency compared to traditional schemes. Compared to Lazy-WL and SWANS, PS-WL decreases the average response time by at least 98.4% and up to 99.1%
Notably, both PS-WL and our ablation variant DegradedWL maintain extremely low and stable response times. This indicates that the core architecture, which leverages a data protection area to offload I/O-intensive wear-leveling operations, is highly effective at minimizing performance overhead. The primary performance gain comes from this architectural design rather than the specific lifetime metric used.
@@
Figure 1: Average Response Time result of different SSD array scaling configuration
4.2.2 Lifetime Standard Deviation
A lower standard deviation of lifetime among disks signifies a more effective wear balance. As shown in Fig.2 , PS-WL achieves a superior wear-balancing effect with exceptional stability. On average, PS-WL outperforms Lazy-WL by 90.5%, SWANS by 88.1%, and, most importantly, our ablation variant DegradedWL by 48.3%.
This superior performance, especially the consistent low deviation of PS-WL, stems from its adaptive parameter tuning mechanism, which allows it to dynamically adjust to different scaling configurations. In contrast, the other schemes, lacking this adaptivity, exhibit significant performance fluctuations across different scaling scenarios. For instance, in Fig. 2(a), the standard deviation for SWANS and Lazy-WL varies dramatically between different (k,k’) configurations, whereas PS-WL maintains a consistently low and predictable level of deviation. This highlights the robustness and adaptability of our approach.
Figure 2: Lifetime Standard Deviation result of different SSD array scaling configuration
4.2.3 Triggered Wear-Leveling Operations
We evaluate the efficiency of each scheme by measuring the total number of pages migrated for wear-leveling. A lower count indicates less overhead. According to Fig.3, the wear-leveling operations triggered by PS-WL are reduced on average by 95.5% and 85.2% compared to SWANS and Lazy-WL, respectively. This demonstrates PS-WL’s exceptional efficiency in achieving wear balance with minimal data movement.
Interestingly, PS-WL triggers slightly more wear-leveling operations than DegradedWL, with an average increase of 11.0%. This is an expected and deliberate trade-off. The additional, more precisely targeted migrations are driven by the probability-sensing model to proactively mitigate risks that the P/E cycle metric cannot perceive. As the next section will show, this marginal increase in overhead is justified by a substantial reduction in array-level failure risk.
Figure 3: Triggered Wear-Leveling Operations result of different SSD array scaling configuration
4.2.4 Analysis of Probability-Sensing
As shown in Fig. 4 (a-d) , PS-WL’s proactive approach yields a significantly lower final array-level failure probability. On average, PS-WL reduces this probability by 61.4% compared to DegradedWL. This demonstrates that by sensing the imminent risk, PS-WL correctly accelerates its wear-leveling process to exit the high-risk state more rapidly. By achieving the wear-leveling completion state faster, PS-WL ensures the array spends a much longer subsequent operational period in a state of low, stable failure probability.
This enhanced reliability and faster risk mitigation is achieved at the cost of a controlled 11.0% increase in wear-leveling operations, as confirmed in Fig. 4 (e-h). This result powerfully demonstrates our core thesis: by sensing failure probability, PS-WL makes more intelligent and timely data migration decisions, trading a small, controlled increase in overhead for a major improvement in both the final array reliability and the speed at which that reliability is achieved.
Figure 4: Analysis of Probability-Sensing
4.3 Analysis
Our evaluation demonstrates that PS-WL substantially outperforms prevailing inter-disk wear-leveling methodologies. Compared to established schemes like SWANS and Lazy-WL, PS-WL slashes average response time by over 98.8% and cuts wear-leveling operations by 85.2% to 95.5%. This efficiency stems from its latency-friendly data conservative zone, which effectively mitigate data migration overhead.
Beyond these performance gains, PS-WL’s primary innovation is its intelligent and adaptive nature. The ablation study against DegradedWL provides clear evidence of this: by replacing the conventional P/E cycle metric with our nuanced failure probability model, PS-WL reduces lifetime standard deviation by an additional 48.3%. This result confirms that sensing actual failure risk leads to far more effective and stable wear-balancing decisions.
This superior balancing is achieved through a deliberate and highly effective trade-off. A controlled 11.0% increase in wear-leveling operations—driven by more precise risk assessments—enables PS-WL to slash the final array failure probability by 61.4% in high-risk scenarios. Ultimately, the results confirm that PS-WL maintains exceptional stability and performance across all tested configurations, validating the robustness of its probability-sensing model and adaptive PID control.
@@ 5 Conclusion
The experimental evaluation validates the effectiveness of our approach. Our results show that PS-WL achieves up to an 88% reduction in lifetime standard deviation, a 98% decrease in average response time, and slashes wear-leveling data migration by up to 95% compared to state-of-the-art methods. Most critically, it reduces the final array failure probability by over 60% in high-risk scenarios when compared to a non-probabilistic control, proving its superior capability for intelligent risk mitigation. Collectively, these findings establish PS-WL as a highly effective and robust solution, well-suited for the demands of modern, scalable SSD storage systems.

### v2

@@ 2.3 Existing Wear Leveling Approaches
  • EDM: Endurance-aware Data Migration[19] prolong SSD endurance and boost system performance by minimizing the write amplification inherent to the migration process. The method incorporates a hot-data-first policy, giving priority to the migration of hot data with the highest write frequency to quickly balance the load. Furthermore, a supplementary cold-data-first policy is used to migrate infrequently accessed cold data.
@@ 3.4 Wear Leveling Control
$\frac{|L_{o}(t)-L_{s}(t)|}{L_{o}(t)}<\lambda$ (8)
where $\lambda$ is a preset exit threshold.
@@ 4.1 Methodology
To evaluate the performance of our proposed PS-WL scheme, we conduct a comprehensive comparison against three baseline approaches: SWANS[8], Lazy-WL[9] and EDM[19], with different scaling methods in various RAID configurations. We used the Alibaba Cloud Block Storage Traces[20], a write-intensive, large-scale I/O trace for our experiments.
@@ 4.2 Experimental Results
We conducted tests on the SSD emulator we designed for $(k_{o},k_{s})$ scaling scenario where $k_{o}$ is the number of original disk and $k_{s}$ is the number of scaling disk. To quantify the direct benefit of our probability-sensing model, we designed a specific experiment comparing PS-WL against other WL approach. .
4.2.1 Lifetime Standard Deviation
The primary objective of this section is to evaluate the effectiveness of the PS-WL scheme in achieving wear balance among drives within an SSD array. We use the standard lifetime deviation as the core metric for assessing array-level wear-leveling effectiveness. A lower lifetime standard deviation indicates that the wear-leveling strategy has effectively distributed the write workload across all disks. To facilitate a cross-scheme comparison, we represent this metric using the standard deviation of the Program/Erase (P/E) cycle counts of all disks in the array.
The experimental results demonstrate that PS-WL consistently maintains the lifetime standard deviation at an extremely low level across all scaling scenarios and workloads, significantly outperforming all comparative schemes. This directly proves the precision and effectiveness of the PS-WL strategy in tracking and rectifying wear imbalances. In contrast, SWANS, Lazy-WL, and EDM produced higher lifetime standard deviations in most cases, indicating that their wear-leveling approaches have significant deficiencies in adaptability across various scenarios.
Figure 1 and 2 present the results for write-intensive workloads. PS-WL reduced the lifetime standard deviation by an average of approximately 79.9%, 88.5%, and 83.7% compared to SWANS, Lazy-WL, and EDM, respectively.
Figure 1: Lifetime Standard Deviation result when adding original disks
Figure 2: Lifetime Standard Deviation result when adding scaling disks
@@ 4.2.2 Average Response Time
This section aims to evaluate the impact of the PS-WL scheme on system I/O performance during SSD array scaling, specifically measured by the key metric of Average Response Time. In enterprise-level storage systems, maintaining a stable and low-latency response time is crucial, particularly when executing background tasks such as data migration and wear-leveling operations. A lower average response time signifies that the scaling process causes less disruption to system performance.
The experimental results clearly demonstrate that the PS-WL scheme exhibits superior performance across all tested combinations. Its average response time is consistently maintained at an extremely low level, and it shows exceptional stability. In contrast, the three comparative schemes—SWANS, Lazy-WL, and EDM—all induced significant performance jitter and increased latency, with their average response times being far higher than that of PS-WL.
As shown in Figure 3 and 4, under write-intensive workload, the effects of write amplification and garbage collection on SSDs become more pronounced. This, in turn, further amplifies the potential impact of wear-leveling operations on performance. Compared to SWANS, Lazy-WL, and EDM, the PS-WL scheme reduced the average response time by approximately 92.8%, 97.5%, and 86.4%, respectively.
Figure 3: Average response time result when adding original disks
@@
Figure 4: Average response time result when adding scaling disks
4.2.3 Wear-Leveling Trigger Count
This section analyzes the Wear-Leveling Trigger Count, a metric that records the total number of times the wear-leveling algorithm was activated to perform data migration throughout the simulation. It directly quantifies the operational overhead of the wear-leveling scheme itself. An efficient algorithm should achieve the desired balancing effect with an appropriate number of triggers and a suitable amount of data migration. Triggering too frequently can introduce unnecessary performance overhead and additional write amplification, whereas triggering too infrequently may indicate that the wear-leveling strategy is sluggish.
Figure 5: Wear-Leveling Trigger Count result when adding original disks
Figure 6: Wear-Leveling Trigger Count result when adding scaling disks
Among these schemes, SWANS lacks a module to restrain wear-leveling operations, leading it to trigger the most wear-leveling actions in many configurations, far exceeding other schemes. The trigger counts for Lazy-WL and EDM exhibit significant fluctuations, causing their wear-leveling control to be overly aggressive in some configurations while appearing inactive in others. On average, PS-WL improved upon SWANS, Lazy-WL, and EDM by 73.9%, 34.4%, and 13.8%, respectively.
The results from the evaluation metrics above lead to the conclusion that PS-WL does not merely pursue the lowest short-term data migration cost. Instead, by executing moderate and highly efficient migration operations, it achieves a degree of wear balance far superior to other schemes, thereby maximizing the long-term reliability and lifespan of the entire SSD array. This effectiveness-oriented, adaptive overhead control strategy is one of the core advantages of the PS-WL scheme over the other baseline approaches.
4.2.4 Total I/O Count
The Total I/O metric records the total number of I/O operations required by different wear-leveling schemes to achieve a state of wear balance among the drives in the array under various scaling scenarios. This metric measures the efficiency of wear-leveling; a lower Total I/O count indicates that a state of wear balance is reached more quickly. To validate the effectiveness of our proposed scheme, the experiment uses different initial wear states and compares the speed at which the exit condition for the array’s wear-leveling mechanism is met. This serves to evaluate PS-WL’s adaptive balancing speed when facing different initial conditions and to verify its contribution to enhancing overall array reliability.
As observed in the preceding comparative experiments, the schemes based on SWANS, Lazy-WL, and EDM fail to converge wear-leveling in all configurations across multiple scenarios. Therefore, comparing the Total I/O count across schemes in a context of unlimited wear-leveling has little reference value. Consequently, to demonstrate PS-WL’s adaptive capability for SSD array scaling under varying initial wear conditions, the evaluation of this metric employs an ablation study comparison scheme. This ablation scheme includes all modules except the effective lifetime model, thereby allowing for an assessment of the effectiveness of the effective lifetime concept.
To simulate the scaling of a long-use SSD array in a real-world scenario, the experiment was set up with the original disks having an average initial failure probability of 0.01%. Using a write-intensive workload, we tested the number of I/O operations required to reach an array-level state of wear balance, as well as the Array Failure Probability at that point. This latter metric evaluates the long-term benefits of the scheme from a reliability perspective and is defined as the probability of any single disk in the array failing, which would represent data loss or the need for data reconstruction.
For algorithms represented by Lazy-WL and our ablation experiment, which use wear count differences as the control model input, their wear-leveling speed control remains unchanged under the same initial conditions of average wear disparity. However, as shown in Figure 7, under high initial wear conditions, the PS-WL scheme, using the effective lifetime model, can respond as the failure probability of SSDs in the array gradually increases. It triggers more wear-leveling operations to complete the wear catch-up process more rapidly.
Figure 7: Observing total I/O counts of SSD array with different initial wear, under different scaling configurations in write-intensive workload
The results reveal a critical difference in balancing efficiency. In terms of Total I/O (the line graph), the ablation scheme (blue line) required significantly more I/O operations to achieve wear balance than the complete PS-WL scheme (red line) across all data redistribution strategies and scaling configurations. This gap was particularly pronounced in certain setups. For instance, under the Fastscale scaling scheme with a configuration of three original disks and one scaling disk, the ablation scheme executed approximately 1.63 million I/Os to meet the wear catch-up exit condition, whereas PS-WL required only about 570,000 I/Os.
The data from the bar chart further corroborates this analysis, presenting a contrast in reliability. The ablation scheme is not only inefficient but also poses a higher reliability risk. Due to the massive volume of additional writes it performs, the overall array failure probability for the ablation scheme (blue bars) is much higher than that of the PS-WL scheme (red bars) upon reaching a balanced state. By leveraging its precise perception of effective lifetime, PS-WL can achieve the wear-balancing goal with the minimum number of migrations and the least wear cost. Its wear-leveling strategy can "protect" disks with increasing failure probabilities, thereby maintaining the entire array’s failure risk at an extremely low level.
In summary, this ablation study proves the effectiveness of the proposed failure-probability-based effective lifetime model. This core innovation enables PS-WL to overcome the limitations of traditional P/E count-based balancing. When handling the scaling of SSD arrays with high initial wear, it simultaneously achieves a lower balancing cost (fewer I/Os) and a higher balancing quality (lower failure probability), maintaining a low array-level failure probability over a longer period and demonstrating clear advantages in both efficiency and reliability.
@@ 5 Conclusion
Comprehensive simulation results demonstrate that the PS-WL scheme exhibits significant advantages across multiple dimensions. Compared to existing schemes, it substantially optimizes the wear-leveling performance and response time of SSD arrays under various scaling configurations and test conditions, effectively extending the overall lifespan of the system. This validates its feasibility and advancement in both theoretical design and practical application.